1.

A 400 mH coil of negligible resistance is connected to an AC circuit in which an effective current of 6 mA is flowing. Find out the voltage across the coil if the frequency is 1000 Hz.

2.

A capacitor of capacitance 102/π µF is connected across a 220 V, 50 Hz A.C. mains. Calculate the capacitive reactance, RMS value of current and write down the equations of voltage and current.

3.

Find the impedance of a series RLC circuit if the inductive reactance, capacitive reactance and resistance are 184 Ω, 144 Ω and 30 Ω respectively. Also calculate the phase angle between voltage and current.
1
Expert's answer
2020-04-10T13:40:03-0400

1) The voltage can be found by Ohm's law:


"V=IX_L=I\omega L=I\cdot2\pi fL=15\text{ V}."

2) The capacitive reactance is


"X_C=\frac{1}{2\pi fC}=98\space\Omega."

RMS value of current:


"I=\frac{V}{X_C}=2.2\text{ A}."


3) The impedance of a series RLC circuit is


"Z=\sqrt{(X_L-X_C)^2+R^2}=50\space\Omega."

The phase angle between voltage and current equals


"\phi=\text{atan}\frac{X_L-X_C}{R}=53^\circ."